If you are travelling at an average speed of 60 km/hr, how long will it take you to cover a distance of 180km?
1
Expert's answer
2020-09-28T08:10:21-0400

We can find the time required to cover that distance from the formula:


"t = \dfrac{s}{v} = \dfrac{180 \ km}{60 \ \dfrac{km}{hr}} = 3 \ hr."

Answer:

"t = 3 \ hr."